# UniGraph: Learning a Cross-Domain Graph Foundation Model From Natural Language

### v1

@@ Appendix B Datasets
Cora (He et al., 2024). The Cora dataset consists of 2708 scientific publications classified into one of seven classes – case based, genetic algorithms, neural networks, probabilistic methods, reinforcement learning, rule learning, and theory. The citation network consists of 5429 links. We collect raw text from  (He et al., 2024).
@@
PubMed (He et al., 2024). The Pubmed dataset consists of 19,717 scientific publications from PubMed database pertaining to diabetes classified into one of three classes – Experimental induced diabetes, Type 1 diabetes, and Type 2 diabetes. As in (Liu et al., 2023a), we ask ChatGPT to generate a detailed description of each category. The citation network consists of 44,338 links. We collect raw text from  (He et al., 2024).
@@ Appendix C Implementation Notes
Table 9: Pre-training hyper-parameters for our framework on ogbn-Papers100M.
@@
Table 10: Graph instruction tuning hyper-parameters for our framework on ogbn-Arxiv and FB15K237.
@@
For Wiki-CS, we follow the official splits (Mernyei & Cangea, 2020) with 20 different training splits, we report the average accuracy on the 20 different training splits with 20 random initialization. In each split, 5% of the nodes in each class are used for training.
@@ Appendix E Extended Related Work
Single graph learning. Graph Neural Networks(GNNs) (Kipf & Welling, 2016a; Veličković et al., 2018a) take node features as input and aggregate local neighbor representations using the message-passing paradigm, directly optimizing for specific downstream tasks, achieving superior performance. With GNNs as the backbone, graph self-supervised learning (Kipf & Welling, 2016b; Veličković et al., 2018b) learns representation extractors on unlabeled graphs, subsequently applying the representations to downstream tasks or fine-tuning the pretrained GNNs. Unlike NLP tasks, graph learning tasks exhibit considerable diversity in their forms, making it a challenge to adapt a graph model to different downstream tasks. Recently, graph prompt learning has attempted to unify all tasks into either edge-level (Liu et al., 2023b) or graph-level (Sun et al., 2023) tasks.
@@
Large language models on graphs. Many real-world graphs naturally come with text as node or edge features, which we refer to as Text-Attributed Graphs(TAGs). To facilitate the learning of graph models, using language models to encode text into low-dimensional vectors is a common practice (Hu et al., 2020). Furthermore, recent Large Language Models(LLMs), represented by ChatGPT, demonstrate extensive common knowledge and capabilities as general task solvers. Some works transform graph structures into text and combine them with task descriptions to form prompts as input for LLMs, attempting to directly use LLMs as predictors for handling graph learning tasks (Guo et al., 2023; Wang et al., 2023; Liu & Wu, 2023). Furthermore, some research attempts to enhance the ability of LLMs to understand graph structures and tasks by employing instruction tuning (Ye et al., 2023; Tang et al., 2023).
Appendix F Few-Shot Transfer Results
In Table 11, we report few-shot transfer results with variance.
Table 11: Experiment results in few-shot transfer. We report accuracy (%) for both datasets. Performance of OFA, which is directly trained on test datasets with labels, is marked in gray. Prodigy is pre-trained on the MAG240M or Wiki. UniGraph and other self-supervised learning baselines are pretrained on ogbn-Papers100M, and then evaluated on individual test dataset.
Arxiv FB15K237
40-way 5-way 3-way 40-way 10-way 5-way
Use word2vec to encode raw text as input features.
word2vec 14.42$\pm$0.34 43.24$\pm$0.76 54.24$\pm$0.82 33.93$\pm$0.47 68.24$\pm$0.24 75.80$\pm$0.28
DGI 15.67$\pm$0.22 46.12$\pm$0.84 57.33$\pm$0.23 31.67$\pm$0.55 67.75$\pm$0.35 74.26$\pm$0.73
BGRL 17.98$\pm$0.56 48.43$\pm$1.17 60.24$\pm$0.33 29.24$\pm$0.24 67.23$\pm$1.65 74.14$\pm$0.45
GraphMAE 19.12$\pm$0.64 49.24$\pm$0.14 62.34$\pm$0.87 32.07$\pm$0.45 69.75$\pm$0.23 77.24$\pm$0.72
GraphMAE2 18.45$\pm$0.45 50.01$\pm$0.72 61.35$\pm$0.45 33.01$\pm$1.12 70.45$\pm$0.24 78.01$\pm$1.63
Use LMs to encode raw text as input features.
Prodigy 25.13$\pm$0.07 61.52$\pm$0.28 73.09$\pm$0.36 59.58$\pm$0.22 81.1$\pm$0.39 88.02$\pm$0.48
OFA 22.13$\pm$0.89 60.12$\pm$0.24 72.17$\pm$0.36 65.23$\pm$1.77 83.01$\pm$1.98 90.11$\pm$0.86
Use raw text as input features.
GIANT-XRT 20.12$\pm$0.33 54.33$\pm$0.48 59.98$\pm$0.67 52.63$\pm$0.86 77.21$\pm$0.56 85.57$\pm$0.78
+GraphMAE2 27.35$\pm$0.33 66.91$\pm$0.84 74.62$\pm$0.96 47.73$\pm$0.81 74.33$\pm$0.86 80.17$\pm$0.81
[…]
UniGraph 31.35$\pm$1.01 74.12$\pm$0.83 83.24$\pm$0.97 68.76$\pm$0.54 85.32$\pm$0.85 91.12$\pm$0.73
@@ Appendix G Instruction Prompts
[…]
@@ Appendix H Analysis of LLMs Choices
Table 12: Analysis of LLMs Choices. We report accuracy (%) for all datasets.
[…]
OFA 24.01 56.92 54.01 - - - - - 18.43 19.98 30.96
GraphGPT - - 70.11 - - - - - - - -
  • 1
@@
Table 13: Results of zero-shot transfer compared to ChatGPT. We report accuracy (%) for all datasets.
[…]
GPT-3.5 Turbo 67.69 - 93.42 74.40 - - - - - - -
  • 1

### v2

@@ Appendix A Datasets
Cora (He et al., 2024a). The Cora dataset consists of 2708 scientific publications classified into one of seven classes – case based, genetic algorithms, neural networks, probabilistic methods, reinforcement learning, rule learning, and theory. The citation network consists of 5429 links. We collect raw text from  (He et al., 2024a).
@@
PubMed (He et al., 2024a). The Pubmed dataset consists of 19,717 scientific publications from PubMed database pertaining to diabetes classified into one of three classes – Experimental induced diabetes, Type 1 diabetes, and Type 2 diabetes. As in (Liu et al., 2023a), we ask ChatGPT to generate a detailed description of each category. The citation network consists of 44,338 links. We collect raw text from  (He et al., 2024a).
@@
Appendix B The Acquisition of Text-Attributed Graphs
TAGs are not a newly emerged, narrowly applied research topic. The benchmark datasets widely used in the field of graph learning (such as Cora, PubMed, OGB, etc.) are inherently TAGs, but researchers have pre-encoded them into vector features using methods like word2vec. This is also a major motivation for us to attempt to train a graph foundation model from TAGs: a large amount of graph data is collected from the web, where text and graphs naturally have a close relationship.
Automatic Textual Description Generation. Using natural language to replace vector encoding for representing graph semantic features, converting non-TAGs to TAGs can be an effective method. Although our model can only be used on TAGs, in the paper we also explored the effects of converting non-TAGs, such as molecular graphs, into TAGs. We further conduct experiments to demonstrate that this method of constructing features is comparable, or even superior, to standard vector features. In Table 11, “Standard” refers to standard 9-dimensional node features, containing atomic number and chirality, as well as other additional atom features such as formal charge and whether the atom is in the ring or not. “Text” represents the corresponding textual features we construct, while “text+word2vec” indicates that we use word2vec to encode these texts into vectors.
Table 11. Performance comparison of different feature construction methods on HIV and PCBA datasets.
Features Method HIV PCBA
Standard GAT 70.89$\pm$1.23 56.11$\pm$1.23
Text+word2vec GAT 71.12$\pm$1.34 56.24$\pm$1.01
Text UniGraph 71.23$\pm$1.93 57.67$\pm$0.85
Similarly, for graph data without features, the traditional approach is to construct vector features from each node’s structural features (such as node degree, etc.). We can also use natural language to describe these semantics, thereby converting them into TAGs.
@@ Appendix C Implementation Notes
Table 12. Pre-training hyper-parameters for our framework on ogbn-Papers100M.
@@
Table 13. Graph instruction tuning hyper-parameters for our framework on ogbn-Arxiv and FB15K237.
@@
Table 14. Analysis of LLMs Choices. We report accuracy (%) for all datasets.
Cora PubMed Products1 Wiki-CS WN18RR
OFA 24.01 56.92 54.01 - - - - - 18.43 19.98 30.96
GraphGPT - - 70.11 - - - - - - - -
  • 1
Table 15. Results of zero-shot transfer compared to ChatGPT. We report accuracy (%) for all datasets.
Cora PubMed Products1 Wiki-CS WN18RR
GPT-3.5 Turbo 67.69 - 93.42 74.40 - - - - - - -
  • 1
@@
For Wiki-CS, we follow the official splits (Mernyei and Cangea, 2020) with 20 different training splits, we report the average accuracy on the 20 different training splits with 20 random initialization. In each split, 5% of the nodes in each class are used for training.
@@ Appendix E Extended Related Work
Single graph learning. Graph Neural Networks(GNNs) (Kipf and Welling, 2016a; Veličković et al., 2018a) take node features as input and aggregate local neighbor representations using the message-passing paradigm, directly optimizing for specific downstream tasks, achieving superior performance. With GNNs as the backbone, graph self-supervised learning (Kipf and Welling, 2016b; Veličković et al., 2018b) learns representation extractors on unlabeled graphs, subsequently applying the representations to downstream tasks or fine-tuning the pretrained GNNs. Unlike NLP tasks, graph learning tasks exhibit considerable diversity in their forms, making it a challenge to adapt a graph model to different downstream tasks. Recently, graph prompt learning has attempted to unify all tasks into either edge-level (Liu et al., 2023b) or graph-level (Sun et al., 2023) tasks.
@@
Large language models on graphs. Many real-world graphs naturally come with text as node or edge features, which we refer to as Text-Attributed Graphs(TAGs). To facilitate the learning of graph models, using language models to encode text into low-dimensional vectors is a common practice (Hu et al., 2020). Furthermore, recent Large Language Models(LLMs), represented by ChatGPT, demonstrate extensive common knowledge and capabilities as general task solvers. Some works transform graph structures into text and combine them with task descriptions to form prompts as input for LLMs, attempting to directly use LLMs as predictors for handling graph learning tasks (Guo et al., 2023; Wang et al., 2023; Liu and Wu, 2023). Furthermore, some research attempts to enhance the ability of LLMs to understand graph structures and tasks by employing instruction tuning (Ye et al., 2023; Tang et al., 2023; Xiong et al., 2024; Yang et al., 2023; Sui et al., 2024).
Technical contributions of UniGraph. Existing graph learning models on TAGs are either GNN-nested Transformers like GraphFormers (Yang et al., 2021) or models like GLEM (Zhao et al., 2022) and TAPE (He et al., 2024a) that decouple LMs and GNNs, optimizing them separately. The primary reasons are due to the large parameter size of LMs and the connectivity of graph structures making the model difficult to scale. Additionally, simultaneously optimizing LMs and GNNs with a significant disparity in parameter size can lead to phenomena such as overfitting. Papers like GLEM (Zhao et al., 2022) have mentioned the difficulties in training such a model. In UniGraph, we utilized subgraph sampling and a self-supervised learning loss function based on MLM to effectively mitigate these issues, making this type of backbone network a viable option.
Appendix F Connections to Graph Out-Of-Distribution (OOD) Generalization
While they share a common ground in aiming to enhance the model’s generalization capabilities, they have distinct objectives, methodologies, and applications.
F.1. Similarities
Both approaches aim to improve the model’s generalization capabilities. Our model seeks to create a robust, pre-trained model that can be adapted to various specific tasks or domains. Similarly, Graph OOD Generalization focuses on ensuring that models perform well on graph data that follow a different distribution from the training set.
F.2. Differences
Objectives: UniGraph, while capable of generalizing across domains, primarily focuses on leveraging large-scale data for comprehensive self-supervised pre-training, with less emphasis on specific distributional shifts. Graph OOD Generalization places a strong emphasis on the model’s ability to handle data from distributions that are not represented in the training set, usually supervised learning.
Model architecture and training: UniGraph is a large, pre-trained model that learns from vast amounts of graph data, capturing a wide variety of patterns and features. The training process involves self-supervised pre-training. In contrast, Graph OOD Generalization strategies may be integrated into various model architectures without the necessity for a large-scale pre-training phase. Techniques such as domain adaptation, invariant feature learning, or meta-learning are often employed to improve OOD performance.
Application scope: UniGraph aims to serve as a versatile foundation for a wide range of downstream tasks and domains by leveraging the power of pre-training. As discussed in our paper, the difficulty lies in the fact that graphs from different domains have different features and label spaces; we choose to use text to unify these spaces. Existing methods for graph OOD (Out-Of-Domain) generalization often focus on the same domain and the same feature/label dimensions, where the training and test sets have inconsistent graph structures and feature distributions (Li et al., 2022; Zhang et al., 2021c). This is primarily represented by transfer learning tasks on molecular graphs (Li et al., 2022). To the best of our knowledge, there is currently no paper focused on addressing the issue of cross-domain graph feature/label space misalignment.
@@ Appendix G Instruction Prompts
Citation networks. Given a citation graph, node represents academic paper with a specific topic. <$\text{node}_{v}$> is featured with its content: [Title], [Abstract]. <$\text{node}_{v}$> and its contextual neighbor nodes {<$\textless\text{node}_{u}\textgreater;u\in\mathcal{V}$} are highly correlated. Question: Which category should <$\text{node}_{v}$> be classified as? Please strictly classify the paper into one of the following categories:[Candidate Labels]. Answer:
Products networks. Given a products graph, node represents a product sold in Amazon with a specific category. <$\text{node}_{v}$> is featured with its content: [Content]. <$\text{node}_{v}$> and its contextual neighbor nodes {<$\textless\text{node}_{u}\textgreater;u\in\mathcal{V}$} are highly correlated. Question: Which category should <$\text{node}_{v}$> be classified as? Please strictly classify the product into one of the following categories:[Candidate Labels]. Answer:
Web networks. Given a Wikipedia graph, node represents Wikipedia page with a specific category. <$\text{node}_{v}$> is featured with its content: [Name],[Content]. <$\text{node}_{v}$> and its contextual neighbor nodes {<$\textless\text{node}_{u}\textgreater;u\in\mathcal{V}$} are highly correlated. Question: Which category should <$\text{node}_{v}$> be classified as? Please strictly classify the Wikipedia page into one of the following categories:[Candidate Labels]. Answer:
Knowledge graphs. Given a knowledge graph, edge between two entities represents a relation with a specific category. Node one <$\text{node}_{v}$> is featured with its content: [Name],[Content]. Node two <$\text{node}_{u}$> is featured with its content: [Name],[Content]. Question: Which category should the relation between node one Node one <$\text{node}_{v}$> and node two <$\text{node}_{u}$> be classified as? Please strictly classify the Wikipedia page into one of the following categories:[Candidate Labels]. Answer:
@@ Appendix H Analysis of LLMs Choices
Appendix J On Incorporating Additional Graph Zero-Shot Learning Baselines
Table 16. Zero-shot results comparison between UniGraph and OpenGraph on Cora and PubMed datasets.
Method Cora PubMed
OpenGraph 75.04 68.69
UniGraph 78.24 74.21
Table 16 compares the zero-shot results of UniGraph and OpenGraph (Xia et al., 2024) on the Cora and PubMed datasets. We adopt an experimental setup strictly consistent with OpenGraph, pre-training on the pre-training datasets provided by OpenGraph.
Table 17. Zero-shot results comparison between UniGraph and LLaGA on Arxiv, Products, Cora and PubMed datasets.
Method Arxiv Products Cora PubMed
LLaGA 74.29 82.21 92.42 87.82
UniGraph 76.01 82.41 93.12 87.98
Table 17 compares the zero-shot results of UniGraph and LLaGA (Chen et al., 2024) on the Arxiv, Products, Cora and PubMed datasets. We adopt an experimental setup strictly consistent with LLaGA, pre-training on the pre-training datasets provided by LLaGA.
Table 18. Zero-shot results comparison between UniGraph and ZeroG on Cora and PubMed datasets.
Method Cora PubMed
ZeroG 68.72 78.02
UniGraph 72.19 81.23
Table 18 compares the zero-shot results of UniGraph and ZeroG (Li et al., 2024) on the Cora and PubMed datasets. We adopt an experimental setup strictly consistent with ZeroG, pre-training on the pre-training datasets provided by ZeroG.